Suppose you have GH¢ 5.00. If you use the money to pay electricity bill, you will gain or benefit, in monetary terms, GH¢ 8.00 whiles if you use it to board a taxi you will gain GH¢ 10.00.

(a) What is the opportunity cost for spending the money on transportation (taxi)? (b) What is the opportunity cost for spending the money on electricity bill?

(c) What is the net benefit when

(i) the direct cost is used?

(ii) the opportunity cost is used?

(d) From your results in (cii) which good should you purchase? Explain your answer.


1
Expert's answer
2022-02-15T13:34:29-0500

(a) What is the opportunity cost for spending the money on transportation (taxi)?


Opportunity cost can also be calculated simply by dividing the cost of what is given up by what is gained.

Thus the opportunity cost for spending the money on transportation is given by:

Opportunity cost = "\frac{cost\, of\, transport}{cost\, of elctricity\,bill}" = "\frac{8}{10}" = 0.8

Thus, the opportunity cost for spending the money on transportation is GH¢. 0.8


b) What is the opportunity cost for spending the money on electricity bills?


Opportunity cost can also be calculated simply by dividing the cost of what is given up by what is gained.

Thus the opportunity cost for spending the money on electricity bills is given by:

Opportunity cost = "\frac{cost\, of\, of elctricity\,bill }{cost\, of\,transport}" = "\frac{10}{8}" = 1.25

Thus, the opportunity cost for spending the money on electricity bills is GH¢. 1.25


(c) What is the net benefit when

(i) the direct cost is used?


Net benefit = total benefit - total cost

= (10 + 8) - (5 + 5)

Thus Net benefit = GH¢ 8


(ii) the opportunity cost is used?


Opportunity Cost of taxi = sum of Explicit Cost (taxi) + Implicit Cost (forgone profits from electricity bill) to Pursue.

= 5 + 5 = 10

Thus, the opportunity cost for spending the money on transportation is GH¢. 8


Opportunity Cost of E. bills = sum of Explicit Cost (E. bill) + Implicit Cost (forgone profits from taxi) to Pursue.

= 5 + 3 = 8

Thus, the opportunity cost for spending the money on electricity bills is GH¢. 8


(d) From your results in (cii) which good should you purchase? Explain your answer.

I will spend money electricity bill.

Because the economic profit benefit of paying electricity bill is 10 - 8 = GH¢. 2